OpenStudy (anonymous):

solve the equation by completeling the square p^2 + 6p + 8 =0

OpenStudy (bahrom7893):

p^2 + 2*3p + 8 + 1 - 1 = 0

OpenStudy (bahrom7893):

p^2 + 2*3p +9 = 1 (p+3)^2 = 1

OpenStudy (bahrom7893):

p+3 = + or -1 p+3 = 1 p+3=-1

OpenStudy (bahrom7893):

p = -2 p = -4

OpenStudy (nenadmatematika):

correct :D